# Formatting Instructions For NeurIPS 2025

### v1

@@ Appendix A Supplemental Implementation Details
Appendix B Supplemental Ablation Study on $\tau$
@@ Appendix D Limitations and Future Work
While our work demonstrates an adequate acceleration of video LLMs by token merging for inference, it is important to outline its current limitations. First, the approach is primarily designed for fixed-length video clips and does not natively support online, arbitrary-length streaming video input. This poses challenges for real-time processing [5, 29, 28] and maintaining long-term context understanding. Second, as shown in Fig. 5, similar to other methods [6, 45, 48, 36] in the token pruning area, our approach does not optimize the latency of the vision tower. Further work, such as quantization [24, 44, 13] and methods to accelerate the vision tower [9, 39, 42], is worth exploring for further optimization.

### v3

@@ B.3 Enhanced Performance with Reduced Overhead
B.4 Supplemental Ablation Study on $\tau$
@@
Table 6: Ablation Study on Merge Strategy. Our mixed strategy achieves the best performance.
Method MVBench EgoSchema LongVideoBench VideoMME Avg. Score
Attention 58.4 60.9 55.9 57.4 58.1
DPC-KNN 57.4 59.6 53.9 56.6 56.9
HoliTom 58.4 60.9 56.2 58.3 58.5
B.5 Ablation Study on Merge Strategy
The core of spatio-temporal merging in HoliTom is a hybrid strategy that integrates attention-guided compression with similarity-guided clustering (DPC-KNN). As shown in Tab. 6, our mixed strategy achieves the best performance. The key to this lies in distinguishing between the two token types encountered during the merging process. For non-redundant tokens (within a single frame), the encoder’s self-attention scores are excellent priors, making an attention-guided compression strategy highly effective. However, for redundant tokens (formed by merging tokens from adjacent frames), the original single-frame self-attention scores lack a theoretical basis as a merging metric. Therefore, DPC-KNN clustering is adopted to group these merged tokens based on feature similarity, which provides a more principled and effective approach in this specific spatio-temporal context.
@@ Appendix D Limitations and Future Work
While our work demonstrates an adequate acceleration of video LLMs by token merging for inference, it is important to outline its current limitations. First, the approach is primarily designed for fixed-length video clips and does not natively support online, arbitrary-length streaming video input. This poses challenges for real-time processing [5, 33, 32] and maintaining long-term context understanding. Second, as shown in Fig. 5, similar to other methods [6, 53, 57, 41] in the token pruning area, our approach does not optimize the latency of the vision tower. Further work, such as quantization [27, 52, 13, 16], methods to accelerate the vision tower [9, 45, 49] and new application areas [50, 46, 54], is worth exploring for further optimization.